OpenStudy (anonymous):

Show steps please If you invest $4000 in an account paying 5.6% compounded continuously, how long will it take for your money to accumulate to $7000?

OpenStudy (slaaibak):

\[7000 = 4000e^{n * 0.056}\] \[{7000\over4000} = e^{n * 0.056}\] ln both sides: \[\ln({7000\over4000}) = \ln(e^{0.056*n})\] Using ln laws:

OpenStudy (mathmate):

If the money is to compound continuously, the formula to use would be: FV=PVe^(rt) 7000=4000e^(0.056)t Take log both sides and divide, solve for t to get t=9.993 years

OpenStudy (slaaibak):

\[\ln({7\over4}) = 0.056n * \ln e = 0.056n\] \[n = { \ln({7\over4}) \over 0.056 }\]

OpenStudy (anonymous):

slaaibak. Why does the equation I used give a different value?

OpenStudy (slaaibak):

Because if you compound something continuously, the formula is A = Pe^nr

OpenStudy (mathmate):

Your equation is for compound interest. The question requires a continuous compounding, which means that the compounding period is infinitely short. It requires the formula involving e.

OpenStudy (anonymous):

ty mathmate. Good to know :-)

OpenStudy (slaaibak):

To observe why this happens, evaluate the following limit: \[\ \lim_{x \rightarrow \infty} (1 + {1\over x})^x\]